OpenStudy (anonymous):

What is the solution to the following equation? x2 + 7x - 3 = 0

OpenStudy (anonymous):

do you know the formula?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Good question !

OpenStudy (anonymous):

\[ x=\frac{-b + \sqrt {b^2-4ac}}{2a}\quad\text{and}\quad x=\frac{-b - \sqrt {b^2-4ac}}{2a} \]

OpenStudy (anonymous):

where Ax^2 + Bx + C = 0

OpenStudy (anonymous):

\[\LARGE x_{1/2}=\frac{-b\pm \sqrt{b^2-4ac}}{2a}\] \[\LARGE ax^2+bx+c=0\] in your case you have: \[\huge \underbrace{1}_{a}\cdot x^2\underbrace{+7}_{b}x \underbrace{-3}_{c}=0\]